OpenStudy (anonymous):

A ball is thrown vertically upward at an initial speed of 128 feet per second. If the distance from the ground t seconds after being thrown is y=128t-16t^2, neglecting air resistance, find: The maximum height that the ball will reach before beginning to fall. At what time does it attain its maximum height? At what time does it fall on the ground? How long is the ball in the air? At what time (s) is the ball 240 feet from the ground?

OpenStudy (tkhunny):

You shoul dhave some background in the Parabola. Let's see your best work. Solve a Quadratic Equation. Find an Axis of Symmetry. Go!

OpenStudy (anonymous):

let y =0 then find the derivative of 128t-16t^2 0=128t-16t^2 0=128-32t 32t=128 t=4sec. the maximum height reached